A machine gun Fires 50gm bullets at a speed of 1000m/s. The gunner, holding the machine gun in his hands, can exert an average force of 180N against the gun. Determine the maximum number of bullets he can fire per minute


1
Expert's answer
2021-10-05T13:41:50-0400

It is the impulsive force so...


"I=mv"


=50×10-3kg×1000=50


If man can hold the gun at a recoiling force of 180N so in 1 min


180∗60=10800


"No. Of bullets per min =\frac{ force * time }{impulse control}"

"= \frac{10800}{50}"

= 216